OpenStudy (anonymous):

@mertsj Given P(E) = 0.58 and P(F n E) = 0.32, find P(F|E)

OpenStudy (mertsj):

\[P(F|E)=\frac{P(EnF)}{P(E)}\]

OpenStudy (anonymous):

I dont know how I would solve it.

OpenStudy (mertsj):

What is P(EnF)?

OpenStudy (anonymous):

0.32

OpenStudy (mertsj):

What is P(E)?

OpenStudy (anonymous):

0.58

OpenStudy (anonymous):

0.55 is the answer?

OpenStudy (mertsj):

What is: \[\frac{P(EnF)}{P(E)}\]

OpenStudy (anonymous):

0.32/0.58 = 0.55

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

Okay thank you

OpenStudy (mertsj):

yw